If \(P(A) = 0.35\), then \(P(A') = \)
A. \(0.35\)
B. \(1.35\)
C. \(0.75\)
D. \(0.65\)
A. \(0.35\)
B. \(1.35\)
C. \(0.75\)
D. \(0.65\)
Best answer
Answer: D. \(0.65\)
\(P(A') = 1 - P(A) = 1 - 0.35 = 0.65\).
Log in or create a free account to answer this question.